Half life period of U234 is 2.5×105 years. In how many years it will remain 25% of original amount?
1
Expert's answer
2020-04-08T10:28:44-0400

"N_0" - initial amount of U234

"N=0.25 \times N_0" - final amount of U234

"T=2.5 \times 10^5 \text{ years}" - half life period

We can use the fact that


"N=N_0 \Big( {1\over 2}\Big)^{\frac{t}{T}}"



"N=0.25\times N_0=N_0 \Big( {1\over 2}\Big)^{\frac{t}{T}} \ \Rightarrow 0.25= \Big( {1\over 2}\Big)^{\frac{t}{T}} , \ \Big( {1\over 2}\Big)^2=\Big( {1\over 2}\Big)^{\frac{t}{T}} \ \Rightarrow 2=\frac{t}{T}"

"t=2T=5.0 \times 10^5 \text{ years}"


Answer: in "5\times 10^5" years it will remain 25% of original amount.